[Hallc_running] [New Logentry] Hall C RC update: 10/24/2021 (no RC meeting today)

mack at jlab.org mack at jlab.org
Sun Oct 24 09:45:01 EDT 2021

Logentry Text:
We've been back in good production at ~42muA since 6:15pm on Saturday. 

I'm a little nervous our good luck is because they've imposed lower beam currents on our Hall A brethren, and the problem of A beam apparently going thru the C slit and causing BLM trips in Arc 7 isn't yet resolved. (I can see them sneaking down the Hall C current even now on the CEBAF Status Information page lol.) 

No RC meeting today. Sorry for the late notice. I will nevertheless update the RC meeting page later today once I have time to put together something more coherent. 

Updates from the 9am PD meeting: The Hall A pass change and Secretary of Energy tour will take place on Monday. Probably the only thing that's going to happen on Monday regarding the Arc 7 magnet short is to inspect the magnet more and take measurements for construction of a drift pipe that will replace the magnet if it is going to be pulled out. So after we recover beam on Monday, expect continued delivery to A,B,C but no Hall D. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3931581
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211024/09471a6d/attachment.html>

More information about the Hallc_running mailing list